OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope.A(5, 5), m = 3

jimthompson5910 (jim_thompson5910):

Start with point slope form y - y1 = m(x - x1) plug in the given slope to get y - y1 = 3(x - x1) then plug in the given point to get y - 5 = 3(x - 5)

jimthompson5910 (jim_thompson5910):

from here, you need to write the equation in Ax+By = C form y - 5 = 3(x - 5) y - 5 = 3x - 3*5 ... distribute y - 5 = 3x - 15 ... multiply y - 5 - 3x = -15 ... Subtract 3x from both sides. y - 3x = -15+5 ... Add 5 to both sides. y - 3x = -10 ... Combine like terms. -3x + y = -10 ... rearrange terms 3x - y = 10 ... multiply every term by -1 to change the sign of the x term